OpenStudy (owlet):

TOPIC: DIFFERENTIAL EQUATIONS Verify that the function is general solution of nonhomogenous differential equation. Function is \(\sf y=c_1•cos(x) + c_2•sin(x)+ x•sin(x) + (cos(x))ln(cos(x))\) DE: \(\sf y"+y=sec\ x\) Interval: \(\sf (-\frac{\pi}{2}, \frac{\pi}{2})\)

OpenStudy (owlet):

from the theory, the general solution is the sum of the solution of the homogenous differential equation \(\sf y_c\) and a particular solution of the nonhomogenous differential equation \(\sf y_p\): \(\sf y=y_c + y_p\) So I am expecting to get: \(\sf y_c=c_1•cos(x) + c_2•sin(x)\) \(\sf y_p=x•sin(x) + (cos(x))ln(cos(x))\) So I solved for \(\sf y_c\) first: \(\sf y"+y=0\\ let\ y=e^{mx} \\ ... \\ m^2+m= 0\\ m(m+1)=0\\m=-1,\ m=0 \\ hence, y_c=c_1e^{-x} + c_2\) which doesn't match with what I am expecting..... (P.S. We haven't covered the Variation of parameters yet, so I can't use that)

OpenStudy (irishboy123):

\(m^2 \color{red}{+ 1}\)

OpenStudy (owlet):

ohhhh right , thanks.